Shift Summary for Day 1-21-00

Subject: Shift Summary for Day 1-21-00
From: Vivian O'Dell (
Date: Fri Feb 04 2000 - 15:35:11 CST

              Shift Summary for day shift, 21/Jan/99

On Shift: E. Turner, A. Bellavance (+ RickK, Ashkan, Theo)

Hours of beam delivered at all intensities: 4.7
Hours of beam delivered at nominal* intensity: 4.2
Hours of EM regenerator data: 3.6

Accounting of lost time: TeV down, access to move BAG, changing
configurations, Kessler tests

Runs ended this shift:
Run 15655 - 5.7M events - BEAM_EM
Run 15656 - 744K events - BEAM_EM (Note: for this run the NM4AN magnet was
                                   in 799 mode instead of 832 mode)

Tapes brought to FCC this shift: none

Accesses during shift:

BobH (with Theo and Albert) to move BAG from E to W, and remove quartz

RickK and Anatoly to move quatrz (but they didn't really do it).

Overview of shift:
  EM regenerator datataking

Description of Shift:

run spill time comment
               7:00 Preparing for EM reg. run. Ramping up NM4AN.
                        20" absorber in. Beam request is 3.5e12.
15655 1 7:20 Start run.
  " 7:30 Request 4e12. Actual SEM is hovering around 3e12.
  " 30 7:51 RickK takes one spill of accidental-only data with
                        the absorber completely out. Takes 1-2 spills to
                        recover (pipeline, KFREND, etc.)
  " 8:00 SEM is finally settling around 4e12.
  " 50 8:20 Kessler takes another accidental-only
                        one-shot. Recover by spill 53.
  " 58 8:30 TeV abort. Beam goes away.
               8:55 Ch.13:"TeV extraction problems; investigating."
  " 71 Turn OFF magnet at NM3S. No beam yet.
  " 77 9:02 Beam returns.
  " Notice NM2S3 has tripped off. Call MCR to reset.
  " 84 Kessler takes another accidental-only one-shot.
  " 110 9:50 Change to 10" absorber.
  " 145 10:28 TeV abort. Beam goes away.
  " 10:35 Ch.13:"TeV link problems; investigating."
  " 150 10:40 Beam returns.
  " 159 10:53 Turn magnet NM3S back ON.
  " 11:56 BobH arrives wanting to go on access. Call MCR.
15655 last 12:00 Beam goes away, stop the run.
              12:05 BobH, Theo, and Albert go on access. Moved BAG
                        from east to west hole and removed quartz counter
                        from btw BAG and UCal. The quartz is on the
                        raised floor next to the laser power supply.
              12:26 Done with access.
              12:29 Beam returns.
15656 1 12:38 Do drt_fresh_start and begin new BEAM_EM run.
  " 13:00 Tev abort and switchyard energy conservation (beam
  " 13:04 Ch.13:"F4 fridge problems; investigating."
15656 13:58 Ch.13:"Replacing F4 turbo cart. Estimate 1 hour
                        to beam." Run stopped, experiment turned over to
                        RickK for beam size test.
              14:41 RickK and Anatoly on access to move quartz
                        counter to behind BAG.
              14:50 Return from access without doing anything.
              15:06 Ch.13:"F4 turbo cart replacement
                        continues. Est. beam return at 16:30."
              15:30 Hand off to evening shift.

Follow up Notes on Previous Shift

Ignored red SDAQ buttons "OnSplit" and "VolMon2 (UCLA stuff)"

Items to Follow up on Next Shift

Ignore red SDAQ buttons "OnSplit" and "VolMon2 (UCLA stuff)"

Angela Bellavance work: (630) 840-2801 home: (630) 548-3412 page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
                                - Arthur C. Clarke

This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:35:14 CST